What is the largest value of X makes this number 982,7X1,588 divisible by 9
Answer by jim_thompson5910(35256)

The digits must add to some number that's a multiple of 9.

9+8+2+7+x+1+5+8+8

x+48

x + 45 + 3

(x+3) + 45

(x+3) + 9*5

So x+3 must be a multiple of 9. If x is a single digit number, then x must be 6 since x+3 = 6+3 = 9

So the largest value of x is 6.